OpenStudy (anonymous):

John invested part of his $6000 savings in common stock and rest in rare stamps. At the end of the year , she realized a gain of 9% on the stock and 12% on the stamps. If his savings now amount to $6615, hoq much did ge invest in stamps

OpenStudy (anonymous):

Without figuring the amounts yet and just using a variable like "x", can you write an expression for how much was invested in common stock and then how much was used for the gain on stamps? This is to get you started and the question is designed for you to get yourself started.

OpenStudy (anonymous):

I didn't get it

OpenStudy (anonymous):

If you have a fixed quantity of something to start with and you take away a variable or unknown amount, how much is the unknown amount taken away and what is the amount left?

OpenStudy (anonymous):

This should get you started. How much do you have to start with?

OpenStudy (anonymous):

$6000

OpenStudy (anonymous):

If you are taking away a variable or unknown amount, how much are you taking away? Hint: use a variable, like a letter.

OpenStudy (anonymous):

x+y=6000?

OpenStudy (anonymous):

That would be true, but this problem uses only one variable. If you take away "x" amount of money for use as an investment vehicle for stamps, then you have "6000 - x" left for use in your investment in stock because: x + (6000 - x) = 6000 No need to introduce a second variable.

OpenStudy (anonymous):

But x+(6000-x)=6000 is equal to 0

OpenStudy (anonymous):

No, what you have is: 6000 = 6000 You already have your equal sign. You already have an equation. You will have 615 in profit because you end up with 6615 at the end, so your interest profit is 615: (0.12)(x) + (6000 - x)(0.09) = 615 So, now just solve for "x".

OpenStudy (anonymous):

X=2500?

OpenStudy (anonymous):

Is that right? 2500?

OpenStudy (anonymous):

Yes! Good job! So, 2500 is used for his stamp investment and the rest is used for stock.

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

More importantly, do you understand now how to set up the equation? That's where the real work is.

OpenStudy (anonymous):

And you're welcome. Nice working with you.

OpenStudy (anonymous):

Yes i know now, thank you !

OpenStudy (anonymous):

Good luck to you in all of your studies and thx for the recognition! @Kittykatkat14